1. Solve the ODE/TVP: y +2y+y=5(1-2),y(0)-0.7(0) =0. Use the Convolution Theorem everywhere possible, in parts (b) and (c).
Answer #1

t O O + dr + t dr = et / dr This can also be written as y(t) = (t-2) et 2 ult-2) Therefore by Coridution property of Laplace